The pH in a 0.13 M solution of aniline (Kb=4.3×10−10). pH=8.87

Calculate the concentration of C6H5NH+3 in a 0.13 M solution of aniline.

Calculate the concentration of OH− in a 0.13 M solution of aniline.

Calculate the concentration of H3O+ in a 0.13 M solution of aniline.
Answer #1

Aniline is C6H5NH2

Kb=4.3x10-10

Dissociation of aniline in solution is as shown

Initial concentration (M) 0.13 0 0

Change -y +y +y

Final concentration 0.13-y y y

Kb=4.3x10-10==(y x y)/(0.13-y)

4.3x10-10=(y x y)/(0.13-y)

As the value of Kb is very small, we can ignore y as compared to 0.13 as y<<<<0.13 M

So 4.3x10-10=(y x y)/(0.13)

4.3x10-10 x 0.13=y2=0.559 x 10-10

y=0.75 x 10-5 M

So =0.75 x 10-5 M

We know that ionic product of water Kw=1.0x10-14=[H3O+][OH-]

Substituting the value of [OH-]

1.0x10-14=[H3O+] x 0.75 x 10-5 M

[H3O+]=1.0x10-14/0.75 x 10-5 M=1.33 x 10-9

pH=-log [H3O+]=-log (1.33 x 10-9)=8.87